OpenStudy (kohai):

The slope of the line tangent to the graph of ln(xy) = x at the point where x=1 is...

OpenStudy (anonymous):

slope is given by dy/dx so diffrentiate ln(xy) with respect to x

OpenStudy (kohai):

This is an extra credit problem because we haven't worked with ln. I'm a quick study, we just haven't learned this yet

OpenStudy (anonymous):

u know how to diffrentiate

OpenStudy (kohai):

I'm not sure how to differentiate ln

OpenStudy (anonymous):

diffrentiation of lnx is 1/x

OpenStudy (anonymous):

here 1/xy(dy/dx)=1

OpenStudy (kohai):

Oh, okay, so the slope is 1?

OpenStudy (kohai):

1/xy(dy/dx)=1 I assumed that that meant the slope is 1

OpenStudy (kohai):

Okay, continue please

OpenStudy (kohai):

There's 0, 1, e, e^2, and 1-e for the options

OpenStudy (kohai):

@perl

OpenStudy (perl):

hello

OpenStudy (anonymous):

@gorv

OpenStudy (kohai):

Perl would you mind explaining how to do this problem? @danish071996 and I have tried a wack at it but are stumped

OpenStudy (perl):

implicit derivative 1/(xy) * (1*y + x* y' ) = 1

OpenStudy (kohai):

That just all went over my head. I'm sure I'll understand shortly, we haven't done ln in class yet, this is an extra credit homework problem

OpenStudy (perl):